In a school, 30 percent of the students do not travel by bus. From a sample

of 10 students find the probability that (a) at least 3 students travel by bus

(b) at most 4 students travel by bus (c) between 2 and 7 clients students

travel by bus (exclusive)



1
Expert's answer
2021-06-25T07:06:46-0400

Let "X=" the number of students who travels by bus: "X\sim Bin (n, p)."

Given "n=10, q=0.3, p=1-q=1-0.3=0.7."

(a)


"P(X\geq3)=1-P(X=0)-P(X=1)-P(X=2)"

"=1-\dbinom{10}{0}(0.7)^{0}(0.3)^{10-0}-\dbinom{10}{1}(0.7)^{1}(0.3)^{10-1}"

"-\dbinom{10}{2}(0.7)^{2}(0.3)^{10-2}=0.9984096136"

"\approx0.9984"

(b)


"P(X\leq4)=P(X=0)+P(X=1)+P(X=2)"

"+P(X=3)+P(X=4)=\dbinom{10}{0}(0.7)^{0}(0.3)^{10-0}"

"+\dbinom{10}{1}(0.7)^{1}(0.3)^{10-1}+\dbinom{10}{2}(0.7)^{2}(0.3)^{10-2}"

"+\dbinom{10}{3}(0.7)^{3}(0.3)^{10-3}+\dbinom{10}{4}(0.7)^{4}(0.3)^{10-4}"

"=0.0473489874\approx0.0473"

(c)


"P(2<X<7)=P(X=3)+P(X=4)"

"+P(X=5)+P(X=6)=\dbinom{10}{3}(0.7)^{3}(0.3)^{10-3}"

"+\dbinom{10}{4}(0.7)^{4}(0.3)^{10-4}+\dbinom{10}{5}(0.7)^{5}(0.3)^{10-5}"

"+\dbinom{10}{6}(0.7)^{6}(0.3)^{10-6}=0.009001692"

"+0.036756909+0.1029193452"

"+0.200120949=0.3487988952\approx0.3488"